OpenStudy (anonymous):

Please help! Simplify: 5+x/x+5

OpenStudy (anonymous):

Is it 1?

OpenStudy (dumbcow):

yes good job

OpenStudy (katrinakaif):

This would simpily = 1 I believe

OpenStudy (anonymous):

Thank you guys :)

OpenStudy (dumbcow):

commutative property...you change the order around 5+x = x+5

OpenStudy (anonymous):

yup! 5+x=x+5, you u can make ur problem 5+x/5+x

myininaya (myininaya):

=1 when x does not equal -5

myininaya (myininaya):

but i think they want you to assume that x+5 is never zero